# Comparing next-generation detector configurations for high-redshift gravitational wave sources with neural posterior estimation

### v1

@@ Appendix C Learning curves and p-p plots
Table 5 also reports the Bayesian evidence $\mathcal{Z}(d)$ defined in Equation D1, along with its uncertainty estimated by averaging the weights (Equation 1) obtained with Dingo-IS (mcbook):
$\mathcal{Z}(d)\pm\sigma=\frac{1}{n}\sum_{i}w_{i}\left(1\pm\sqrt{\frac{1-\epsilon}{n-\epsilon}}\right),$ (D5)
@@
where $n$ denotes the total number of samples drawn from the proposal distribution $q_{\phi}(\theta|d)$, and $\epsilon$ is the sampling efficiency defined in Equation 2. The log evidences computed with Dingo-IS and Bilby are consistent within statistical uncertainties.
@@ Appendix E Information gain
We compute Equation 3 through a Monte Carlo integration:
$I=\int p(\theta|d)\log_{2}\frac{p(\theta|d)}{\pi(\theta)}\mathrm{d}\theta\approx\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\log_{2}\frac{p(\theta_{i}|d)}{\pi(\theta_{i})},$ (E1)
where $N_{s}=3\times 10^{4}$ denotes the number of samples drawn from the posterior. In our case, these samples are obtained through sampling–importance–resampling (SIR; Rubin1988UsingTS) based on the importance weights defined in Equation 1. Using Bayes’ theorem, we can rewrite the ratio inside the logarithm as:
$I=\frac{1}{N_{s}}\sum_{i}^{N_{s}}\log_{2}\frac{p(\theta_{i}|d)}{\pi(\theta_{i})}=\frac{1}{N_{s}}\sum_{i}^{N_{s}}\log_{2}\frac{\mathcal{L}(d|\theta_{i})\cancel{\pi(\theta)}}{\mathcal{Z}(d)\cancel{\pi(\theta)}}=-\log_{2}\mathcal{Z}(d)+\frac{1}{N_{s}}\sum_{i}^{N_{s}}\log_{2}\mathcal{L}(d|\theta_{i}),$ (E2)
where the log evidence $\log_{2}\mathcal{Z}(d)$ is evaluated through the importance sampling weights (Equation D5). We estimate the variance associated with the Monte Carlo estimator of the information gain (2003itil.book.....M; Talbot:2023pex; Heinzel:2025ogf) as follows:
$\sigma_{I}^{2}=\frac{1}{N_{s}}\left[\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(-\log_{2}\mathcal{Z}(d)+\log_{2}\mathcal{L}(d|\theta_{i})\right)^{2}-I^{2}\right],$ (E3)
where $I$ is defined in Equation E2 and $\sigma_{I}^{2}<10^{-3}$ bits for all detector configurations.
@@ Appendix F Multimodal luminosity distance
We apply Hartigan’s dip test for unimodality (fb099965-2cf5-3955-b5e5-077ac9290f24; 4c99dc66-50c0-3bf4-8c15-3b2f3d6ec1eb) to the one-dimensional marginal posterior of the luminosity distance. A distribution is classified as unimodal if the $p$-value exceeds 0.01. As shown in Figure 10, about 19% (23%) of sources have a multimodal $D_{\mathrm{L}}$ posterior for 2L MisA (2L A), while the $\Delta$ configurations display only $\sim 2$%.
Appendix G Sky modes in the 2L Misaligned configuration
@@
Figure 11: Marginalized one- and two-dimensional posterior distributions for all parameters, with the exception that this time we show the source-frame chirp mass ($\mathcal{M}_{\mathrm{s}}$) and redshift ($z$). The results for the low-mass event (orange) are shown alongside those for the high-mass event (blue). Vertical and horizontal lines indicate the true injected values and contours represent the 68% and 95% credible regions. Top-right panel: Sky map of the low-mass (orange) and high-mass (blue) events, with contours indicating the 50% and 90% credible area. See Appendix G for further details.
@@
Figure 2 illustrates multimodalities in sky localization for a source observed with the 2L MisA configuration. The presence of these multimodalities is further confirmed in Figure 4, where almost no sources exhibit a single mode. Most sources ($\gtrsim 60\%$) display between two and six sky modes, while $\sim 20\%$ exhibit eight distinct modes.
@@
In this section, we provide an intuitive example to illustrate that the presence of multiple sky modes in the 2L MisA configuration is specific to the sources considered in this study, after applying the sample efficiency selection (Section III.1 and bottom panel of Figure 1). For a small subset of sources that are excluded from our analysis but still present in the prior space, it is reasonable to expect that the number of sky modes would be significantly lower. Intuitively, the baseline between the 2L ET interferometers becomes effectively long for sources with sufficiently high SNR above a certain frequency threshold, which depends on both the length of baseline and the source’s sky location. A comprehensive study to quantify this effect is beyond the scope of this manuscript and will be addressed in future work.
@@
Figure 11 shows the one- and two-dimensional marginal distributions obtained with Bilby for two sources at zero noise: a less massive, closer source with $\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}=55.5~\mathrm{M}_{\odot}$ and $D^{\mathrm{inj}}_{\mathrm{L}}=6.2~\mathrm{Gpc}$, with optimal SNR = 87.5, and a more massive, more distant source with $\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}=145.3~\mathrm{M}_{\odot}$ and $D^{\mathrm{inj}}_{\mathrm{L}}=37.2~\mathrm{Gpc}$ with optimal SNR = 28.7. The sky location, source-frame chirp mass, mass ratio, and other parameters are fixed for both sources to the maximum-likelihood values estimated for GW250114 (LIGOScientific:2025rid), while only the luminosity distance—and thus the redshift—are varied.
Figure 11 illustrates that multimodalities in sky localization vanish for the less massive source, whereas they persist for the more massive one, yielding $\Delta\Omega_{90\%}=3$ deg2 for the low-mass event and $\Delta\Omega_{90\%}=194$ deg2 for the high-mass event.
Appendix H Performance in estimating other parameters
Figure 12: Cumulative distributions of events as a function of the optimal SNR ($\rho$, Equation D4), the relative variations in detector-frame chirp mass ($\Delta\mathcal{M}_{\mathrm{d}}/\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}$), mass ratio ($q$), first ($\Delta\chi_{1}$) and second ($\Delta\chi_{2}$) aligned spins, and inclination angle ($\Delta\theta_{\mathrm{jn}}$) for all considered XG detector configurations. See Appendix H for details.
For completeness, Figure 12 presents the performance of the considered detector configurations in estimating the remaining parameters, such as optimal SNR, detector-frame chirp mass, mass ratio, the aligned spin components, and inclination angle. The overall trend is confirmed, as the $\Delta$ and 2L A configurations provide the poorest performance, while the 2L MisA + CE setup delivers the most accurate estimates..

### v2

@@ Appendix D Dingo-IS versus Bilby
Table 5 also reports the Bayesian evidence $\mathcal{Z}(d)$ defined in Eq. 8, along with its uncertainty estimated by averaging the weights (Eq. 1) obtained with Dingo-IS (Owen, 2013):
$\mathcal{Z}(d)\pm\sigma=\frac{1}{n}\sum_{i}w_{i}\left(1\pm\sqrt{\frac{1-\epsilon}{n-\epsilon}}\right),$ (12)
@@
where $n$ denotes the total number of samples drawn from the proposal distribution $q_{\phi}(\theta|d)$, and $\epsilon$ is the sampling efficiency defined in Eq. 2. The log evidences computed with Dingo-IS and Bilby are consistent within statistical uncertainties.
@@ Appendix E Information gain
We computed Eq. 3 through a Monte Carlo integration:
$I=\int p(\theta|d)\log_{2}\frac{p(\theta|d)}{\pi(\theta)}\mathrm{d}\theta\approx\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\log_{2}\frac{p(\theta_{i}|d)}{\pi(\theta_{i})},$ (13)
where $N_{s}=3\times 10^{4}$ denotes the number of samples drawn from the posterior. In our case, these samples were obtained through sampling-importance resampling (Rubin, 1988) based on the importance weights defined in Eq. 1. Using Bayes’ theorem, we can rewrite the ratio inside the logarithm as
$I=\frac{1}{N_{s}}\sum_{i}^{N_{s}}\log_{2}\frac{p(\theta_{i}|d)}{\pi(\theta_{i})}=\frac{1}{N_{s}}\sum_{i}^{N_{s}}\log_{2}\frac{\mathcal{L}(d|\theta_{i})\cancel{\pi(\theta)}}{\mathcal{Z}(d)\cancel{\pi(\theta)}}=$ (14)
$-\log_{2}\mathcal{Z}(d)+\frac{1}{N_{s}}\sum_{i}^{N_{s}}\log_{2}\mathcal{L}(d|\theta_{i}),$ (15)
where the log evidence $\log_{2}\mathcal{Z}(d)$ is evaluated through the importance sampling weights (Eq. 12). We estimated the variance associated with the Monte Carlo estimator of the information gain (Mackay, 2003; Talbot and Golomb, 2023; Heinzel and Vitale, 2025) as follows:
$\sigma_{I}^{2}=\frac{1}{N_{s}}\left[\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(-\log_{2}\mathcal{Z}(d)+\log_{2}\mathcal{L}(d|\theta_{i})\right)^{2}-I^{2}\right],$ (16)
where $I$ is defined in Eq. 14 and $\sigma_{I}^{2}<10^{-3}$ bits for all detector configurations.
@@ Appendix F Multimodal luminosity distance
We apply Hartigan’s dip test for unimodality (Hartigan and Hartigan, 1985; Hartigan, 1985) to the one-dimensional marginal posterior of the luminosity distance. A distribution is classified as unimodal if the $p$-value exceeds 0.01. As shown in Fig. 11, about 19% (23%) of sources have a multimodal $D_{\mathrm{L}}$ posterior for 2L MisA (2L A), while the $\Delta$ configurations display only $\sim 2$%.
Appendix G Robustness to waveform approximant variations
@@
Figure 12: Marginalized one- and two-dimensional posterior distributions for the detector-frame chirp mass ($\mathcal{M}_{\mathrm{d}}$), mass ratio ($q$), luminosity distance ($D_{\mathrm{L}}$), right ascension (ra), and declination (dec) of the event displayed in Fig. 2, observed with the 2L MisA configuration in zero noise. Results from Bilby are presented for different waveform approximants: XPHM (orange), XO4a (grey), NRSur7dq4 (teal), and SEOBNRv5PHM (blue). Vertical and horizontal lines indicate the true injected values (see Table 5) and the contours denote the 95% credible regions. Further details are provided in Appendix G.
To assess whether the multimodalities in luminosity distance and sky localization observed in Fig. 2 for the 2L MisA configuration persists under different waveform approximants, we repeat the analysis in zero noise using Bilby (see Appendix D). In addition to the fiducial IMRPhenomXPHM waveform model (XPHM; Colleoni et al. 2025; see also Appendix B), we both injected and recovered the signals using three additional state-of-the-art inspiral–merger–ringdown (IMR) approximants: NRSur7dq4 (Varma et al., 2019), SEOBNRv5PHM (Ramos-Buades et al., 2023), and IMRPhenomXO4a (Thompson et al., 2024, XO4a).
All considered waveform models describe precessing quasi-circular binaries and include higher-order multipole contributions; however, they adopt distinct modeling strategies (Chatziioannou et al., 2024). In brief, NRSur7dq4 is constructed by interpolating numerical-relativity simulations (Field et al., 2014; Blackman et al., 2015) and is therefore typically the most accurate for high-mass systems, such as GW231123 (Abac and others, 2025a; Bini et al., 2026). In contrast, XPHM, SEOBNRv5PHM, and XO4a combine analytical frameworks with numerical-relativity calibration to build complete IMR models that are applicable across the full range of total masses (Buonanno and Damour, 1999, 2000; Buonanno et al., 2007; Ajith and others, 2011). The differing modeling strategies and treatments of precession dynamics render these waveform models largely independent. As shown in Fig. 12, we find no evidence of waveform-induced systematics in the inferred parameters, with all four waveform models reproducing the same number and locations of multimodalities.
Appendix H Sky modes in the 2L misaligned configuration
Figure 13: Marginalized one- and two-dimensional posterior distributions for all parameters, with the exception that this time we show the source-frame chirp mass ($\mathcal{M}_{\mathrm{s}}$) and redshift ($z$). The results for the low-mass event (orange) are shown alongside those for the high-mass event (blue). Vertical and horizontal lines indicate the true injected values and contours represent the 68% and 95% credible regions. Top-right panel: Sky map of the low-mass (orange) and high-mass (blue) events, with contours indicating the 50% and 90% credible area. See Appendix H for further details.
@@ Appendix H Sky modes in the 2L misaligned configuration
Figure 2 illustrates multimodalities in sky localization for a source observed with the 2L MisA configuration. The presence of these multimodalities is further confirmed in Fig. 4, where almost no sources exhibit a single mode. Most sources ($\gtrsim 60\%$) display between two and six sky modes, while $\sim 20\%$ exhibit eight distinct modes.
@@
In this section we provide an intuitive example to illustrate that the presence of multiple sky modes in the 2L MisA configuration is specific to the sources considered in this study, after applying the sample efficiency selection (Sect. 3.1 and the bottom panel of Fig. 1). For a small subset of sources that are excluded from our analysis but still present in the prior space, it is reasonable to expect that the number of sky modes would be significantly lower. Intuitively, the baseline between the 2L ET interferometers becomes effectively long for sources with sufficiently high S/N above a certain frequency threshold, which depends on both the length of baseline and the source’s sky location. A comprehensive study to quantify this effect is beyond the scope of this manuscript and will be addressed in future work.
@@
Figure 13 shows the one- and two-dimensional marginal distributions obtained with Bilby for two sources at zero noise: a less massive, closer source with $\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}=55.5\penalty 10000\ \mathrm{M}_{\odot}$ and $D^{\mathrm{inj}}_{\mathrm{L}}=6.2\penalty 10000\ \mathrm{Gpc}$, with optimal S/N = 87.5, and a more massive, more distant source with $\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}=145.3\penalty 10000\ \mathrm{M}_{\odot}$ and $D^{\mathrm{inj}}_{\mathrm{L}}=37.2\penalty 10000\ \mathrm{Gpc}$ with optimal S/N = 28.7. The sky location, source-frame chirp mass, mass ratio, and other parameters are fixed for both sources to the maximum-likelihood values estimated for GW250114 (Abac and others, 2025b), while only the luminosity distance—and thus the redshift—are varied. Figure 13 illustrates that multimodalities in sky localization vanish for the less massive source, whereas they persist for the more massive one, yielding $\Delta\Omega_{90\%}=3$ deg2 for the low-mass event and $\Delta\Omega_{90\%}=194$ deg2 for the high-mass event.